OpenStudy (anonymous):

A recursive formula for a geometric sequence is {g1 = 54 {g_n + 1 = (2/3)g_n for integers n ≥ 1 Write an explicit formula for the sequence.

OpenStudy (anonymous):

g_n=((2/3)^n-1)g_1

OpenStudy (anonymous):

\[ g_n = 54\cdot \left(\frac{2}{3}\right)^{n-1} \]